# Large Language Models Are Active Critics in NLG Evaluation

### v1

@@ 6.2 Explainability Analysis
Average 98.78 94.74 96.05 94.03 - - -
@@
Table 3: Human evaluation of explanations generated by Active-Critic on SummEval samples. It assesses (1) the quality of individual explanations, (2) the alignment of the overall explanation with criteria-specific explanations, (3) the distinguishability of the overall explanation across cases of varying quality, and (4) the overall usefulness of the generated explanations per testing case.
To conduct a deeper analysis of the explanations generated by Active-Critic, we employed three proficient English-speaking annotators to evaluate the quality of the scoring explanations on a random sample of 150 test cases from SummEval. Our assessment consisted of four parts, details in Appendix F. First, for each individual explanation per case, each annotator rated the quality based on: (1) clarity of the statement, (2) relevance to the target criterion, (3) alignment with the corresponding score, and (4) accuracy within the context of the test case (e.g., correctness in matching the source text). Further emphasizing the overall scoring explanation per case, we asked annotators to assess its alignment with the criteria-specific explanations, and its differentiability across cases of varying quality, respectively. Finally, we asked annotators to provide an overall rating on a scale of 1-5 based on the usefulness of all generated explanations per case. To validate the reliability of human annotations, following prior work  (Fabbri et al., 2021), we calculated intercoder reliability by Krippendorff’s alpha (Krippendorff, 2011). The 0.6534 Kappa coefficient indicates substantial agreement among annotators.
@@
Table 3 shows the results. Overall, we observe a comparatively high quality of individual explanations over four considered dimensions, with 98.78% clarity, 94.74% relevance, 96.05% score consistency, and 94.03% information accuracy on average. Most testing cases have overall explanations that align with the criteria-specific ones (95.11%), and 90% of the overall explanations effectively differentiate case quality. With an average rating of $\sim$4.5 out of 5 on the generated explanations across sampled testing cases, the result shows that explanations generated by Active-Critic are of good quality and useful to explain the resulting scores.
Figure 3: Impact of prompt optimization for scoring and mini-batch iterations for task inference. We compare Active-Critic’s performance using Kendall-Tau(%). The Pearson and Spearman results are in Appendix E.
6.3 Ablation Study
Impact of Optimization.
We assess the impact of optimization on Active-Critic by comparing its performance when dynamic prompt optimization for scoring is removed and, additionally, when mini-batch iterations are eliminated during task inference. Figure 3 displays the results. Across all four NLG tasks, we observe a consistent performance drop when prompt optimization for scoring is removed, with a further decline when only using a single mini-batch of labeled data for task inference. This suggests that both scoring prompt optimization and mini-batch iterations for task inference are crucial for Active-Critic to achieve more human-aligned evaluations. Interestingly, the Active-Critic shows greater sensitivity to scoring optimization in the fine-level evaluation of SummEval and the coarse-level evaluation of SFRES, indicating that this component plays a more significant role in these specific evaluation scenarios. In contrast, the influence of mini-batch iterations for task inference is minimal in SummEval, suggesting that Active-Critic can effectively infer the target evaluation task in this setting with limited training data.
SummEval TopicalChat SFRES OpenMEVA (ROC) Average
$\gamma$ $\rho$ $\tau$ $\gamma$ $\rho$ $\tau$ $\gamma$ $\rho$ $\tau$ $\gamma$ $\rho$ $\tau$
Ours (AC-Fine) 0.6301 0.5486 0.4299 0.6023 0.6214 0.4713 0.2915 0.2501 0.1906 0.5259 0.5363 0.4109 0.4591
w/o Task Description 0.5825 0.4826 0.3552 0.4949 0.5057 0.4211 0.2011 0.1406 0.1129 0.3846 0.3802 0.2918 0.3628
w/o Criteria Definition 0.5726 0.522 0.4062 0.5533 0.5368 0.4451 0.2567 0.243 0.1571 0.4176 0.4237 0.326 0.405
w/o McS-E 0.5386 0.5227 0.4156 0.611 0.6173 0.4845 0.3094 0.2663 0.199 0.4908 0.4962 0.3622 0.4428
w/o OS-E 0.6106 0.5129 0.3908 0.5639 0.5615 0.4464 0.2844 0.2113 0.1602 0.509 0.4931 0.3632 0.4256
Table 4: Ablation study of key modules in Active-Critic.
Key Module Analysis.
We further analyze the individual contribution of each module in Active-Critic by comparing performance with and without that module. Table 4 shows the results across four NLG tasks. Noted that, the variant w/o criteria inference uses the original predefined criteria from each benchmark for further computation. In the variant w/o OS-E, we calculated the overall quality score per test case by averaging the multiple criteria-specific scores generated from McS-E.
Overall, we find that removing task inference modules leads to a more substantial performance drop in our protocol compared to scoring modules, especially when the LLM is not asked to infer the task description (resulting in a $\sim$10decrease on average). This suggests that our approach of guiding LLM in determining what to evaluate is more important to Active-Critic’s effectiveness than other modules. The greater performance drop for the variant w/o OS-E, compared to the one w/o McS-E, indicates that the LLM-generated overall quality score contributes more meaningfully than simply averaging the generated criteria-specific scores.
7 Conclusion and Discussion
@@
We proposed Active-Critic, a novel LLM-based NLG evaluation protocol that relies solely on lightweight human-scored data. Unlike existing machine-based evaluators that depend on human-predefined task descriptions and evaluation criteria, Active-Critic actively infers the necessary details about the target evaluation task and provides explainable assessments by drawing insights directly from the data. This paradigm shift will open the door to endow Active-Critic with the adaptability to evaluate systems on new NLG tasks without extensive task-specific prompt engineering. Experiments across four distinct NLG tasks demonstrate LLMs’ potential as active critics, achieving a higher correlation with human judgments compared to state-of-the-art baseline evaluators. Fine-level criteria-specific scoring, paired with the explanation generation setting, prompts the LLM to engage more deeply with the test cases, leading to improved overall quality scoring.
@@ A.3 Instruction component of the evaluation protocol
“Coherence”: “The degree to which the summary flows logically and cohesively, with clear connections between the main points.” “Conciseness”: “The ability of the summary to convey all necessary information in a succinct and efficient manner.” “Coverage”: “The extent to which the summary captures the main events and details from the source text without omitting crucial information.” “Accuracy”: “The faithfulness of the summary to accurately reflect the main points and details of the source text.” “Fluency”: “The readability and naturalness of the language used in the summary, with smooth transitions between ideas and paragraphs.” “Relevance”: “The relevance of the summary to the main topic and the inclusion of only pertinent information from the source text.” “Clarity”: “The clarity and comprehensibility of the summary, with clear and precise language used to convey the main points.” “Engagement”: “The ability of the summary to captivate and engage the reader, drawing them into the main events and details effectively.”
@@ A.4 In-context exemplar of the evaluation protocol
Source: “Paul Merson has restarted his row with Andros Townsend… Any bad feeling between the pair seemed to have passed but Merson was unable to resist having another dig at Townsend after Tottenham drew at Turf Moor.” Response: “Paul merson has restarted his row with andros townsend .. in the 83rd minutefor tottenham as they drew 0-0 against burnley.” Multiple Evaluation Criteria: Coherence: The degree to which the summaryflows logically and cohesively, with clearconnections between the main point. Conciseness: The ability of the summary to convey all necessaryinformation in a succinctand efficient manner. Score Of Each Criterion In JSON: Coherence: {
@@
Response: Chelsea have made an offer for fc … in muto is not connected to the 200million sponsorship deal they signed with japanese company yokohama rubber in February. ”Score of Overall”: 91.66666666666666
@@ A.5 Prompt Template
\${Task Description} Your task is to rate the overall quality of the response, based on the source and the scores for different criteria of the response on a continuous scale from 0 to 100, where 0 means ‘completely irrelevant and unclear’ and 100 means ‘perfectly relevant, clear, and engaging.’ IMPORTANT!! Only output the score as an ‘int’ and nothing else. “Also explain your process to get this score to response. Also please perform error Analysis of given response. What should we change to have a better result?”
@@ Appendix B Details of Parameter Setting and Implementation
Our experiments were carried out using two NVIDIA V100 GPU cards. For prompt optimization in the scoring stage, we utilized the “BootstrapFewShotWithRandomSearch” method in DSPy (Khattab et al., 2023) as the optimizer, which leverages random search to generate examples.
Appendix C Qualitative Analysis of LLM-inferred Criteria.
To analyze the process of automatic evaluation criteria inference by Active-Critic in depth, we conducted a qualitative analysis by reading through three types of criteria per task: (1) criteria generated by AC-Fine, (2) criteria generated by AC-Fine w/o task description, and (3) criteria predefined by humans in each original benchmark. Table 13 shows an illustrative example of SummEval. Moving beyond four primary aspects considered by humans in summarization, Active-Critic also incorporates a range of nuanced criteria, such as clarity, conciseness, coverage/completeness, and engagement. Additionally, each criterion is clearly defined to specify its distinct characteristics.
Coherence: The degree to which the summary flows logically and cohesively, with clear connections between the main points. Conciseness: The ability of the summary to convey all necessary information in a succinct and efficient manner. Coverage: The extent to which the summary captures the main events and details from the source text without omitting crucial information. Accuracy: The faithfulness of the summary to accurately reflect the main points and details of the source text. Fluency: The readability and naturalness of the language used in the summary, with smooth transitions between ideas and paragraphs. Relevance: The relevance of the summary to the main topic and the inclusion of only pertinent information from the source text. Clarity: The clarity and comprehensibility of the summary, with clear and precise language used to convey the main points. Engagement: The ability of the summary to captivate and engage the reader, drawing them into the main events and details effectively. Relevance: How well the response captures the key points or main events from the source text. Conciseness: ”The extent to which the response is brief and to the point, focusing on essential information. Factual Accuracy: Ensuring that the response maintains accuracy and does not introduce new or incorrect information. Clarity: The clarity of language used in the response, avoiding ambiguity or confusion. Structure: The organization and coherence of the response, including the presence of an introductory sentence and logical flow. Completeness: Whether the response covers all the main highlights from the source text without omitting crucial information. Coherence: The overall coherence and logical progression of ideas in the response. Engagement: The ability of the response to engage the reader and maintain interest. Novelty: The extent to which the response introduces new insights or perspectives beyond the source text. Consistency: Ensuring that the responses exhibit a consistent style, tone, and level of detail throughout the dataset. Coherence: the summary should be well-structured and well-organized. The summary should not just be a heap of related information, but should build from sentence to sentence to a coherent body of information about a topic. Consistency: the factual alignment between the summary and the summarized source. A factually consistent summary contains only statements that are entailed by the source document. Fluency: the summary should have no formatting problems, capitalization errors or obviously ungrammatical sentences (e.g., fragments, missing components) that make the text difficult to read. Relevance: The summary should include only important information from the source document.
(a) AC-Fine (b) AC-Fine w/o Task Description (c) Human
Table 13: An illustrative example of the generated evaluation criteria on SummEval, either generated by an Active-Critic variant (a & b) or predefined by humans (c).
Appendix D Impact of Explanations by Pearson and Spearman Correlation
Figure 4: Effectiveness of Explanation in Pearson ($\gamma$).
Figure 5: Effectiveness of Explanation in Spearman ($\rho$). We report the average on Spearman ($\rho$) correlation coefficients for our two variants: AC-Coarse, and AC-Fine, each presented with and without explanation.
@@ Appendix E Impact of Optimization by Pearson and Spearman Correlation
Appendix F Details of Human Evaluation

### v2

@@ 6.3 Qualitative Analysis of Active-Critic
Average 98.78 94.74 96.05 94.03
Table 4: Human evaluation of criterion-specific explanations on SummEval samples.
Dimension Aspect-to-Overall Alignment Differentiability Usefulness
[…]
Overall 95.11 90 4.515
@@
Table 5: Human evaluation of overall explanations on SummEval samples, emphasizing (1) the alignment of the overall explanation with criterion-specific ones, (2) explanations’ differentiability across vary-quality cases, and (3) explanations’ overall usefulness per case.
Human Evaluation of Explanations.
@@
We also employ three proficient English-speaking annotators to evaluate the quality of the scoring explanations generated by Active-Critic on a random sample of 150 test cases from SummEval (see details in Appendix G). As shown in Table 4, the individual explanations demonstrate comparatively high quality across four dimensions, with average scores of 98.78% for clarity, 94.74% for relevance, 96.05% for score consistency, and 94.03% for information accuracy. As shown in Table 5, the overall explanations generally align with the criteria-specific ones (95.11%), and 90% of the overall explanations effectively differentiate case quality. With an average rating of $\sim$4.5 out of 5 on the generated explanations across sampled testing cases, the result shows that explanations generated by Active-Critic are of good quality and useful to explain the resulting scores.
7 Conclusion
@@
We proposed Active-Critic, a novel LLM-based NLG evaluation protocol that relies solely on lightweight human-scored data. Unlike existing machine-based evaluators that depend on human-predefined task-related information for assessment, Active-Critic self-identifies the target evaluation task and nuanced evaluation criteria purely from the data for making judgments. This paradigm shift will enhance the adaptability of Active-Critic, enabling it to flexibly capture the varying priority expectations of different end-users across diverse generation scenarios. Our approach reduces the need for intensive manual efforts to design task-specific criteria and extensive prompt engineering. Experiments across four distinct NLG tasks demonstrate LLMs’ potential as active critics, achieving higher correlation with human judgments compared to baselines. Fine-level criteria-specific scoring, paired with explanations, prompts the LLM to engage more deeply with the test cases, leading to improved overall quality scoring.
8 Limitation
@@ A.3 Instruction component of the evaluation protocol
“Coherence": “The degree to which the summary flows logically and cohesively, with clear connections between the main points." “Conciseness": “The ability of the summary to convey all necessary information in a succinct and efficient manner." “Coverage": “The extent to which the summary captures the main events and details from the source text without omitting crucial information." “Accuracy": “The faithfulness of the summary to accurately reflect the main points and details of the source text." “Fluency": “The readability and naturalness of the language used in the summary, with smooth transitions between ideas and paragraphs." “Relevance": “The relevance of the summary to the main topic and the inclusion of only pertinent information from the source text." “Clarity": “The clarity and comprehensibility of the summary, with clear and precise language used to convey the main points." “Engagement": “The ability of the summary to captivate and engage the reader, drawing them into the main events and details effectively."
@@ A.4 In-context exemplar of the evaluation protocol
Source: “Paul Merson has restarted his row with Andros Townsend… Any bad feeling between the pair seemed to have passed but Merson was unable to resist having another dig at Townsend after Tottenham drew at Turf Moor." Response: “Paul merson has restarted his row with andros townsend .. in the 83rd minutefor tottenham as they drew 0-0 against burnley." Multiple Evaluation Criteria: Coherence: The degree to which the summaryflows logically and cohesively, with clearconnections between the main point. Conciseness: The ability of the summary to convey all necessaryinformation in a succinctand efficient manner. Score Of Each Criterion In JSON: Coherence: {
@@
Response: Chelsea have made an offer for fc … in muto is not connected to the 200million sponsorship deal they signed with japanese company yokohama rubber in February. "Score of Overall": 91.66666666666666
@@ A.5 Prompt Template
\${Task Description} Your task is to rate the overall quality of the response, based on the source and the scores for different criteria of the response on a continuous scale from 0 to 100, where 0 means ‘completely irrelevant and unclear’ and 100 means ‘perfectly relevant, clear, and engaging.’ IMPORTANT!! Only output the score as an ‘int’ and nothing else. “Also explain your process to get this score to response. Also please perform error Analysis of given response. What should we change to have a better result?"
@@ Appendix B Details of Parameter Setting and Implementation
Our experiments were carried out using two NVIDIA V100 GPU cards. For prompt optimization in the scoring stage, we utilized the “BootstrapFewShotWithRandomSearch" method in DSPy (Khattab et al., 2023) as the optimizer, which leverages random search to generate examples.
Appendix C Generalization to Unseen Datasets
Ideally, we expect the Active-Critic-generated evaluation prompts can be directly used for NLG system assessment in a similar future NLG scenario. To assess the generalizability of these prompts, we use the prompts generated by Active-Critic based on SummEval examples to assess unseen cases in Newsroom (Grusky et al., 2018). This dataset comprises 60 news articles and their corresponding summaries generated by 7 summarization systems. Each summary is paired with an overall quality score provided by human annotators. Table 14 displays the results. Our Active-Critic noticeably outperforms baselines by $\sim$10correlation on average, indicating Active-Critic’s generalizability.
Method $\gamma$ $\rho$ $\tau$ AVE
TIGERScore 0.3731 0.41 0.3075 0.3635
UniEval 0.4485 0.4505 0.325 0.408
G-eval (gpt3.5) 0.3853 0.4053 0.3012 0.3639
GPT-3.5 (zero-shot) 0.504 0.561 0.430 0.4983
AC-Fine (GPT3.5) 0.6382 0.6444 0.4949 0.5925
GPT-4 (zero-shot) 0.6583 0.6649 0.4957 0.6063
AC-Fine (GPT4) 0.7466 0.7111 0.5474 0.6684
Table 14: Generalization results of Active-Critic on Unseen Datasets.
Appendix D Additional Results of Active-Critic’s Dependence on Human-scored Data
Figure 4: Results of Active-Critic’s dependence on human-scored data by Pearson, Spearman, and Kendell-Tau, respectively.
@@ Appendix E Impact of Optimization by Pearson and Spearman Correlation
Appendix F Helpfulness of Explanations to Active-Critic’s Judgments
To assess the impact of explanations generated by Active-Critic, we compared our protocol’s performance with versus without explanations, at both coarse and fine levels of evaluations. Figure 7 shows the results based on the Kendall-Tau correlation. We also provide the results of Pearson and Spearman correlation in Figure 8 and Figure 9 respectively.
Figure 8: Effectiveness of Explanation in Pearson ($\gamma$).
Figure 9: Effectiveness of Explanation in Spearman ($\rho$).
Appendix G Details of Human Evaluation Design on Active-Critic’s Explanations
Our assessment consisted of four parts, with details provided below. First, for each individual explanation per case, each annotator rated the quality based on: (1) clarity of the statement, (2) relevance to the target criterion, (3) alignment with the corresponding score, and (4) accuracy within the context of the test case (e.g., correctness in matching the source text). Further emphasizing the overall scoring explanation per case, we asked annotators to assess its alignment with the criteria-specific explanations, and its differentiability across cases of varying quality, respectively. Finally, we asked annotators to provide an overall rating on a scale of 1-5 based on the usefulness of all generated explanations per case. To validate the reliability of human annotations, following prior work  (Fabbri et al., 2021), we calculated intercoder reliability by Krippendorff’s alpha (Krippendorff, 2011). The 0.6534 Kappa coefficient indicates substantial agreement among annotators.